Four electrons are situated at the corners of a rectangle (l=4m; w=2m). Find the resultant force electron A will experience dur to the charges at the other corners of the triangle. The green circle below represents electron A.


1
Expert's answer
2022-02-18T15:53:14-0500

"F_x=kq^2(\frac 1{l^2}+\frac l{(l^2+w^2)^\frac 32}),"

"F_y=kq^2(\frac 1{w^2}+\frac w{(l^2+w^2)^\frac 32}),"

"F=\sqrt{F_x^2+F_y^2}=67.4~\cdot 10^{-30}~N."